OpenStudy (anonymous):

what is the range of f(x)= -2x +5 with domain {1,2,3,4}

OpenStudy (shamil98):

Evaluate at f(1) , f(2), f(3), and f(4) to get the corresponding range values.

OpenStudy (lena772):

range is the difference between the solution to the functions of the highest and lowest value of x

OpenStudy (shamil98):

The range is the value of y , to find it in this case all you do is evaluate when x = 1, x =2, x= 3, x= 4. So when x =1 f(1) = -2(1) +5 = -2 + 5 = 3 so the range is 3 when x = 1, (1,3)